A normal distribution of scores has a variance of S squared = 100. Find the z-scores corresponding to each of the following values:

a) a score that is 20 points above the mean

b) a score that is 10 points below the mean

2) Given a mean score of 43 and a standard deviation of 5 for teaching readiness test among a sample of students calculate raw scores for the following:

a) a z-score of 1.50

b) a z-score of -2.00

3) Subjective wellbeing was measured among a sample of Statistics students with M=150 and S squared=25. Determine the z-scores for the students who obtained the following scores on the subjective wellbeing measure.

a) 110

b) 135





1
Expert's answer
2021-05-31T00:13:56-0400

1. "\sigma = \sqrt{100}=10"

a) "Z = \frac{20}{10} = 2"

b) "Z = \frac{-10}{10}=-1"


2. "\mu=43"

"\sigma = 5"

a) "1.5 = \frac{x-43}{5}"

7.5 = x-43

x=50.5

b) "-2.0=\frac{x-43}{5}"

-10=x- 43

x=33


3. "\mu=150"

"s^2=25"

s=5

a) "Z = \frac{110-150}{5}=-8"

b) Z "= \frac{135-150}{5}=-3"